OpenStudy (1645323):

Which equation is a step in solving 2x2 - 8x + 3 = 0 by completing the square. Select one: a. 2(x2 - 4x + 4) = -3 + 8 b. 2(x2 - 8x + 16) = - 3 + 32 c. 2(x2 - 4x + 2) = -3 + 4 d. 2(x2 - 4x + 4) = -3 + 4

OpenStudy (1645323):

@SolomonZelman

OpenStudy (1645323):

is it A or C?

OpenStudy (solomonzelman):

I never thought of these problems this way. I learned ths in a standard way if you know what I mean. I know that I would have t factor out of 2 to make a perf. sq. on a left side. (x^2-4x+4) would be my perfect square e.i. (b/f factoring out of 2, it would be , 2x^2-8x+8 would do I need to add/subtract from both sides to get 2x^2-8x+8 on the left side? I would be adding 5. SO 2x^2 - 8x + 3 = 0 2x^2 - 8x + 3 + 5 = 0 + 5 2x^2 - 8x + 8= 5 2(x^2-4x+4)=5 so I would say that all of your choices are WRONG, b/c where do you get the 4 from ?

OpenStudy (solomonzelman):

most accurate one is D, but D is also WRONG WRONG and WRONG -:(